# Hybrid Querying Over Relational Databases and Large Language Models

### v1

@@ 4.3. Limitations of HQDL
We believe that the above SQL query offers more control for the database to optimize the hybrid query, build materialized views, and potentially reduce the amount of data generated by LLMs.
5. Evaluation
Table 2. HQDL zero shot evaluation results on the SWAN with the state-of-the art models.
Model California Schools Super Hero Formula One European Football Overall
GPT-3.5 Turbo 50.0% 13.3% 16.7% 16.7% 24.2%
GPT-4 Turbo 50.0% 23.3% 36.7% 16.7% 31.6%
Table 3. HQDL few shot evaluation results on the SWAN benchmark. The numbers in brackets report the accuracy improvement compared to the zero shot method.
Model Fewshot California Schools Super Hero Formula One European Football Overall
GPT-3.5 Turbo 1-shot 50.0% 23.3% 46.7% 26.7% 36.7% (+12.5%)
@@ 5. Evaluation
Table 4. The average F1 score for measuring the facuality of the generated data using HQDL.
Model Fewshot California Schools Super Hero Formula One European Football Average
GPT-3.5 Turbo 0-shot 9.0% 40.5% 19.7% 14.6% 20.9%
1-shot 14.1% 41.7% 54.9% 38.7% 37.3%
3-shot 17.2% 42.4% 66.8% 39.4% 41.4%
5-shot 12.1% 43.0% 66.8% 49.0% 42.7%
GPT-4 Turbo 0-shot 23.0% 43.4% 28.7% 22.1% 29.3%
1-shot 23.7% 45.6% 74.6% 44.3% 47.0%
3-shot 24.1% 45.5% 75.4% 43.4% 47.1%
5-shot 24.0% 45.6% 75.2% 48.1% 48.2%
@@ 5.1. Evaluation Metrics
Data Factuality. We use exact string match to verify the data factuality for each data cell value. Because of the one-to-many relationships (the key from a table map to many values generated by LLMs), we use the widely accepted F1 score, which is a harmonic mean of precision and recall, to measure the overall factuality of generate data entries for each database (Wang et al., 2023b).
Input/Output Tokens. We report the number of tokens (i.e., words, sub-words) used in HQDL which determines the cost.
@@ 5.2. Experiment Configurations
We evaluate HQDL on several OpenAI models (i.e., GPT-3.5 Turbo and GPT-4 Turbo) with the Batch API. In all requests, we set the temperature to 0 and the maximum tokens to 3000.
Few Shots Demonstrations. In the few-shot prompts, we provide static examples randomly selected from the original database. In addition, we added a prefix before the examples to indicate the number of examples provided before the target data entry (i.e., /* k examples are provided before the target data entry */).
5.3. Hybrid Queries Execution Accuracy
@@
In this subsection, we report and analyze the EX scores.
@@
Zero Shot. As shown in Table LABEL:tab:zeroEX, GPT-4 Turbo achieves 31.6% accuracy on the proposed SWAN benchmark, surpassing GPT-3.5 Turbo by 7.4%. One of the main challenges in generating data entries with zero-shot prompts is the data extraction process. Despite specifying the number of fields to be returned in a single row, LLMs sometimes provide too few or too many fields and may occasionally include one or two question marks unchanged in the output. Moreover, we can observe from Table LABEL:tab:zeroEX that for the California Schools and the European Football databases, the GPT-4 Turbo model achieved the same execution accuracy as the GPT-3.5 Turbo model. Since many queries contain a limit clause (users often only want to retrieve the top results), generating more accurate content for irrelevant entries does not necessarily lead to improvements in query execution accuracy. This observation motivated us to further examine the data factuality F1 scores.
@@
Few Shots. The distinction between the few-shots and zero-shot experiments is the inclusion of several static examples. We know that the capabilities of language models can be increased with examples provided for in-context learning (Brown et al., 2020). Hence, we expect LLMs to generate more accurate data. Hence, the execution accuracy should improve as more examples are provided. As shown in Table 3, in general the execution accuracy improves for both models as more examples are provided. When the prompt contains 5 static examples, GPT-3.5 Turbo achieves 38.3% and GPT-4 Turbo achieves 40.0% execution accuracy, 14.1% and 8.4% accuracy improvements compared to the zero shot method. It is also interesting to note that both models achieve around 50% execution accuracy on the California Schools and Formula One database. Upon examining the beyond-database questions for these two databases, we find that one-third of the queries contain a LIMIT clause, retrieving the top or bottom 1 to 10 records. In contrast, only one-tenth of the queries for the Super Hero database contain a LIMIT clause. This suggests that selecting a subset of records using the LIMIT clause plays a significant role in the execution accuracy of hybrid queries. One explanation for this phenomenon is that LLMs may exhibit biases, as previous research has shown that they tend to favor higher socioeconomic entities (Manvi et al., 2024). For instance, while LLMs can accurately identify the champion of a Formula One race, they may struggle to identify the slowest racer and are even more likely to fail in identifying all racers in the same event.
5.4. Data Factuality
@@
To measure data factuality, we use the exact string match to compare the generated data with the ground truth for each cell. Also, we compute the average F1 score over all cells for each database. When the generated content is identical to the ground truth, then it scores a 100% F1 score. As shown in Table 4, GPT-4 Turbo consistently generates more factual information than GPT-3.5 Turbo using the same prompt. For instance, with the 5-shot prompt, GPT-4 Turbo scores 5.5% higher than GPT-3.5 Turbo. Also, the results clearly showcase that providing more examples in the input prompt increases the factuality of the generated output, which also leads to higher execution accuracy when executing the hybrid queries.
5.5. Costs of HQDL
@@
HQDL generated all data entries for the dropped columns. Consequently, for the zero-shot prompt, a total of 6.3 million tokens are used as inputs for LLMs, and 1.5 million tokens are generated by LLMs. On average, about 52k input tokens and 12k output tokens are used per hybrid query. If the number of beyond-database questions increases, the cost per question will decrease. As discussed in Section 4.3, HQDL has room for improvements. We believe that optimizing the pipeline for executing hybrid queries can significantly reduce the token usage and hence reduce the overall costs. Currently, we are using the OpenAI batch APIs which provide a 50% reduction in costs. Hence, to further reduce the price, we believe that a promising direction is to: i) invoke LLMs within SQL and ii) support asynchronous execution for hybrid queries.
6. Conclusion
In this paper, we present the first benchmark, SWAN, for evaluating hybrid queries that answer beyond-database questions using relational databases and large language models. In addition to the benchmark, we also introduce HQDL, a preliminary solution for addressing questions in SWAN. Our evaluation showcase that HQDL can achieve an accuracy of 40.0% in execution accuracy and a 48.2% F1 score for data factuality. We envision that our benchmark along with the HQDL discussions will spark interests within the community to develop comprehensive data systems that leverage the full potential of relational databases and large language models.

### v2

@@ 4.2. Hybrid Query UDFs
Hybrid querying through UDFs offer more control for the database to optimize the hybrid query, build materialized views, and potentially reduce the amount of data generated by LLMs.
Since all four databases utilize SQLite, we can directly leverage BlendSQL (Park et al., 2023), an extended version of the SQLite relational database management system that supports LLM functions. In SWAN, we provide 120 hybrid queries using the BlendSQL syntax, enabling SWAN to evaluate current systems in querying both relational databases and LLMs.
Table 2. HQDL Execution Accuracy results on the SWAN benchmark using different number of demonstrations. The numbers in brackets report the accuracy improvement compared to the zero shot method.
Model Demonstrations California Schools Super Hero Formula One European Football Overall
GPT-3.5 Turbo 0-shot 50.0% 13.3% 16.7% 16.7% 24.2%
1-shot 50.0% 23.3% 46.7% 26.7% 36.7% (+12.5%)
@@
4.3. Optimization Opportunities
While these two solutions (HQDL and Hybrid Query UDFs) can be used to solve the challenges presented in our proposed SWAN benchmark, we believe that there are opportunities to improve upon these two solutions.
First, to answer these beyond database questions, what contexts, other than the necessary keys and the predefined value lists, should be presented in the prompt to reduce LLMs hallucination? There are other attributes inside the relational database that may be relevant and it remains an open question on how to select the best context. One possible approach is to build a vector index on the database values or rows and then fetch the relevant information based on embedding similarity (Zhang et al., 2023; Edge et al., 2024). Second, the prompts and static examples used in HQDL and Hybrid Query UDFs are hand-crafted. It would be more convenient for users if the data system may automatically generate prompts and examples based on the specific context and query requirements. A promising direction is to develop a principled declarative prompt engineering toolkit (Parameswaran et al., 2023). HQDL requires LLMs to generate and materialize all missing data, while Hybrid Query UDFs, through BlendSQL, optimize queries by pushing down predicates to avoid generating unnecessary data entries. Additionally, reusing previously generated data in HQDL is straightforward. In BlendSQL, generated data is cached by mapping the LLM input prompt to its output data. However, prompts with similar meanings (e.g., ”Is the superhero from the Marvel Universe?” versus ”Does the hero come from Marvel?”) cannot directly reuse previous results. A promising approach to address this is incorporating query rewriting within Hybrid Query UDFs to fully leverage all cached LLM-generated data (Zhao et al., 2023). Query optimization and caching are essential for reducing costs and increasing throughput, making hybrid queries more accessible and efficient. BlendSQL currently implemented batching—retrieving data values for multiple rows in a single LLM call—and plans to support parallelized LLM calls in the future to further minimize query latency.
5. Evaluation
Table 3. HQ UDFs evaluation results on the SWAN benchmark. The numbers in brackets report the accuracy improvement compared to the zero shot method.
Model Demonstrations California Schools Super Hero Formula One European Football Overall
GPT-3.5 Turbo 0-shot 10.0% 23.3% 30.0% 10.0% 18.3%
5-shot 13.3% 23.3% 43.3% 3.3% 20.8% (+2.5%)
@@ 5.1. Evaluation Metrics
Table 4. The average F1 score for measuring the factuality of the generated data using HQDL.
Model Demonstrations Average
GPT-3.5 Turbo 0-shot 20.9%
1-shot 37.3%
3-shot 41.4%
5-shot 42.7%
GPT-4 Turbo 0-shot 29.3%
1-shot 47.0%
3-shot 47.1%
5-shot 48.2%
Data Factuality. We use exact string match to verify the data factuality for each data cell value. Because of the one-to-many relationships (the key from a table maps to many values generated by LLMs), we use the widely accepted F1 score, which is a harmonic mean of precision and recall, to measure the overall factuality of generate data entries for each database (Wang et al., 2023b).
Input and Output Tokens. We report the number of input and output tokens (i.e., words, sub-words) used in HQDL and Hybrid Query UDFs, which determine the monetary cost. For instance, GPT 3.5 Turbo priced at \$3 per million input tokens and \$6 per million output tokens.
@@ 5.2. Experiment Configurations
We evaluate HQDL and Hybrid Query UDFs on several OpenAI models (i.e., GPT-3.5 Turbo and GPT-4 Turbo) via OpenAI api calls. In all requests, we set the temperature to 0.
Few Shots Demonstrations. In the few-shot prompts, we provide static examples randomly selected from the original database. For HQDL, the few shots demonstrations are organized as static rows. In Hybrid Query UDFs, the few shots demonstrations are organized as a natural language question, an example database key, and the answer to the natural language question on the example database key (e.g., question: What is the driver code, key: Lewis Hamilton, and answer: HAM).
5.3. HQDL Results
@@
In this subsection, we report and analyze the EX scores and the generated data factuality using F1 score.
@@
Zero Shot. As shown in Table 2, in terms of overall accuracy over all databases using 0-Shot demonstrations, GPT-4 Turbo achieves 31.6% accuracy on the proposed SWAN benchmark, surpassing GPT-3.5 Turbo by 7.4%. One major challenge in using zero-shot prompts to generate data entries lies in ensuring that the output format is consistent, as this significantly impacts the ease of data extraction. Despite specifying the number of fields need to be returned in the input prompt, LLMs sometimes return too few or too many fields and may occasionally return an empty string for a field. Moreover, we can observe from Table 2 that for the California Schools database, GPT-4 Turbo achieved the same execution accuracy as GPT-3.5 Turbo. Questions in the California Schools frequently ask for the top schools. Consequently, queries answering these questions often include a LIMIT clause to retrieve only the top results. Hence, generating more accurate content for irrelevant entries does not necessarily lead to improvements in query execution accuracy. This observation motivated us to further examine the data factuality using F1 scores.
@@
Few Shot. The distinction between the few-shots and zero-shot experiments is the inclusion of several static examples. We know that the capabilities of language models can be increased with examples provided for in-context learning (Brown et al., 2020). Hence, we expect LLMs to generate more accurate data with few shots, and the execution accuracy should improve as more examples are provided. As shown in Table 2, in general the execution accuracy improves for both models as more examples are provided. When the prompt contains 5 static examples, GPT-3.5 Turbo achieves 38.3% and GPT-4 Turbo achieves 40.0% execution accuracy, 14.1% and 8.4% accuracy improvements compared to the zero shot method.
It is also interesting to note that both models achieve the highest execution accuracy on the California Schools and also the lowest execution accuracy on Super Hero database. One-third of the queries in California Schools database contain a LIMIT clause, retrieving the top schools. In contrast, many questions in the Super Hero database seek specific superheroes (e.g., heroes from Marvel or those with blue eyes), and only about one-tenth of the queries for this database include a LIMIT clause. One explanation for execution accuracy difference between questions in California Schools databse and questions in Super Hero database is that LLMs may exhibit biases, as previous research has shown that they tend to favor higher socioeconomic entities (Manvi et al., 2024). For instance, while LLMs can accurately identify schools with the highest standardized testing scores, they may struggle to identify schools with average or below-average grades. Because many queries in the California Schools database contain a LIMIT clause, even when an LLM provides inaccurate answers for many schools, the top results may still appear correct, masking potential errors in the model’s full response.
@@
Data Factuality To measure data factuality (using the F1 score), we use exact string matching to compare the generated data with the ground truth for each cell. Also, we compute the average F1 score over all cells for each database. When the generated content is identical to the ground truth, then it scores a 100% F1 score. As shown in Table 4, GPT-4 Turbo consistently generates more factual information than GPT-3.5 Turbo using the same prompt. For instance, with the 5-shot prompt, GPT-4 Turbo scores 5.5% higher than GPT-3.5 Turbo. Also, the results clearly showcase that providing more examples in the input prompt increases the factuality of the generated output, which also leads to higher execution accuracy when executing the hybrid queries.
5.4. Hybrid Query UDFs Results
We evaluated the performance of BlendSQL (Park et al., 2023) on the SWAN dataset to assess its effectiveness with hybrid query UDFs. Notably, on GPT-3.5 Turbo, the execution accuracy for 0-shot and 5-shot settings reached 18.3% and 20.8% (see Table 3), which are lower compared to HQDL’s results of 24.2% and 38.3%. In our evaluation of hybrid query UDFs, we provided the keys and instructed the LLM to predict only the necessary information (most of the time a single cell value). This approach contrasts significantly with HQDL, where the LLM is given the key and tasked with predicting all column values for the corresponding row. Predicting all column values may be more advantageous than predicting a single column value, as it mirrors a chain-of-thought process that enables the model to leverage inter-dependencies between columns, thereby enhancing accuracy and coherence in its predictions. In HQDL, each LLM call generates a single row. In contrast, BlendSQL uses a default batch size of 5, where each request combines five keys into a list, prompting the LLM to return a list of five data entries corresponding to the five keys. Although batching reduce the number of LLM calls, it also increases the potential for errors, as processing multiple entries in a single call may lead to inaccuracies in the returned data (Cheng et al., 2023).
Another noteworthy difference between HQDL and HQ UDFS is in the format of the few-shot examples. In HQDL, we included static rows in the prompt as few-shot demonstrations, and the goal is to showcase the model how to complete a row of data based on the given keys. In contrast, for HQ UDFS (e.g., BlendSQL), we curated a list of question-answer pairs for each databases, and then BlendSQL selects relevant examples based on similarity metrics (e.g., cosine similarity using a sentence transformer) to find the most similar questions. For example, a demonstration question in the HQ UDFS system might be: ’Provide the city name based on the address.’ Along with this question, given the address ’5328 Brann Street’, the expected city name is ’Oakland’.
5.5. Evaluation Costs
The monetary costs and the system’s performance (e.g., latency and throughout) are implicitly determined by the number of input and output tokens. Here, we report the total number of input and output tokens for HQDL and HQ UDFs.
Table 5. Total tokens used for HQDL and HQ UDFs for zero shot experiments.
Algorithm Input Tokens Output Tokens
HQDL 6.3 M 1.5 M
HQ UDFs 23 M 2 M
@@ 5.5. Evaluation Costs
HQDL generates data entries for all the missing columns. As shown in Table 5, using zero-shot prompt, a total of 6.3 million tokens are used as inputs for LLMs, and 1.5 million tokens are generated by LLMs. On average, about 52k input tokens and 12k output tokens are used per beyond-database questions. If the number of beyond-database questions increases, the cost per question will decrease.
For HQ UDFS, we expected it to use less tokens compared to HQDL, because HQ UDFS give more control to the database query optimizer. This allows the system to intelligently minimize token usage by pushing down predicates, meaning it generates tokens only for the specific data cells needed to answer the query. Surprisingly, for zero-shot prompt, the total input and output tokens are 23 million and 2 million respectively, as shown in Table 5. Compared to HQDL, HQ UDFs uses 3.6x more input tokens and 1.3x more output tokens.
The increased costs of HQ UDFs can be attributed to its limited use of cached results. For instance, to answer the beyond-database question, ’What is the height of the tallest player?’, HQ UDFS used the LLMs to generate heights for all players, as the database lacks this information. Later, another question asks, ’Please list player names who are taller than 180cm.’ The corresponding hybrid query created for this question prompts the LLMs to answer ’Is the player taller than 180cm?’ However, it is evident that the previously generated heights could be directly reused to answer this question, rather than generating new responses. In HQ UDFS, LLM-generated content is cached as a mapping from input prompts to LLM output answers, making it challenging for the system to efficiently reuse cached outputs. In contrast, HQDL stores LLM-generated outputs directly as entities within relationships (schema expansion), simplifying reuse for users.
6. Conclusion and Future Work
In this paper, we present the first benchmark, SWAN, for evaluating hybrid queries that answer beyond-database questions using relational databases and large language models. In addition to the benchmark, we also introduce HQDL, a preliminary solution for answering questions in SWAN based on schema expansion. We also provide queries to evaluate current Hybrid Query UDFs systems (e.g., BlendSQL). Our evaluation demonstrates that there are still many opportunities for improving the execution accuracy and also increasing the overall efficiency. To improve the execution accuracy and ensure high data fidelity, retrieve-augmented generation (RAG) (Li et al., 2024b) and supervised fine-tuning (Elaraby et al., 2023; Razumovskaia et al., 2024) are two promising direction to be integrated in hybrid querying systems. Moreover, there are numerous opportunities to optimize the pipeline for executing hybrid queries, increasing throughput and lowering monetary costs. For example, to further reduce costs and improve system throughput, it is essential to focus on: (i) implementing asynchronous and parallel hybrid query execution, (ii) designing improved caching mechanisms, and (iii) fully utilizing cached content.
Additionally, in the current benchmark, we provided the missing columns for schema expansion and pre-written queries for both HQDL and HQ UDFS. In future work, the process of answering beyond-database questions should be fully automated. Given a natural language question, LLMs should first evaluate whether it can be answered using the existing schema. For questions requiring information beyond the current database, LLMs could be designed to automatically expand the schema, populate missing values, and generate a SQL query (similar to HQDL) or construct a SQL query with user-defined functions to directly prompt LLMs for required information in real time.
We envision that our benchmark, the two baseline solutions, and the discussions on future optimization opportunities will spark interests within the community to develop comprehensive data systems that leverage the full potential of relational databases and large language models.
Acknowledgements.
We thank the anonymous reviewers for their valuable feedback and Parker Glenn for assistance with running BlendSQL on SWAN. Fuheng Zhao was partially funded by Microsoft PhD Fellowship.